# Capacity-Aware Inference: Mitigating the Straggler Effect in Mixture of Experts

### v3

@@ Abstract
The Mixture of Experts (MoE) is an effective architecture for scaling large language models by leveraging sparse expert activation to balance performance and efficiency. However, under expert parallelism, MoE suffers from inference inefficiencies due to imbalanced token-to-expert assignment, where underloaded experts complete computations early but must wait for overloaded experts, leading to global delays. We define this phenomenon as the Straggler Effect, as the most burdened experts dictate the overall inference latency. To address this, we first propose Capacity-Aware Token Drop, which enforces expert capacity limits by discarding excess tokens from overloaded experts, effectively reducing load imbalance with minimal performance impact (e.g., speedup with only degradation on OLMoE). Next, given the presence of low-load experts remaining well below the capacity threshold, we introduce Capacity-Aware Expanded Drop, which allows tokens to include additional local experts in their candidate set before enforcing strict local capacity constraints, thereby improving load balance and enhancing the utilization of underused experts. Extensive experiments on both language and multimodal MoE models demonstrate the effectiveness of our approach, yielding substantial gains in expert utilization, model performance, and inference efficiency, e.g., applying Expanded Drop to Mixtral-87B-Instruct yields a 0.2% average performance improvement and a 1.85 inference speedup.
@@ 1 Introduction
In recent years, the rapid evolution of Large Language Models (LLMs) [32, 38, 11] has driven a wave of innovations, continuously expanding the frontiers of AI research and applications. Among the model architectural innovations, the Mixture of Experts (MoE) framework has emerged as a pivotal technique for optimizing the cost-performance trade-off in LLMs. Specifically, MoE [35, 14] enhances scalability by integrating multiple experts while activating only a subset per input. This selective activation substantially improves model performance without a corresponding increase in computational cost, effectively balancing efficiency and performance.
Despite the success of MoE, a key efficiency challenge lies in the imbalanced token-to-expert distribution, which results in some experts being overloaded while others remain underutilized [26, 46]. In distributed GPU settings, experts are typically sharded across multiple devices, with each GPU responsible for a subset of the experts. Under expert parallelism, low-load experts complete their computations earlier but must wait for overloaded experts to finish, as synchronization barriers are required before proceeding to the next stage. This expert-level straggler effect further propagates to device-level delays, where GPUs hosting lighter expert workloads are stalled by GPUs hosting heavier workloads, leading to inefficient resource utilization and increased end-to-end latency during inference. As illustrated in Figure 1, this phenomenon is referred to as the Straggler Effect, where the heavily loaded experts determine the overall latency of imbalanced MoE inference.
While auxiliary balance losses have been incorporated into the training process to alleviate imbalance [35, 14, 11], these techniques remain ineffective in mitigating imbalance during inference. Specifically, as shown in Figure 1, our findings reveal a highly uneven token distribution among experts, with the highest-load expert handling more than seven times the expected average load. Moreover, managing such imbalance during inference often incurs additional resource overhead. For example, DeepSeek-V3 mitigates this issue by duplicating high-load experts and deploying them redundantly across devices [11]. This motivates us to explore efficient token-to-expert assignment by addressing the key question: How can we prevent extreme overloading of heavily utilized experts?
@@
We identify the Straggler Effect caused by token imbalance at inference time in Mixture of Experts, highlighting the optimization potential for reducing latency.
@@ 2 Related Works
The Mixture of Experts (MoE) is a kind of neural network architecture with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation [21, 23]. The potential of sparse activation in MoE is subsequently exploited by [34] for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision [33] and language [25, 13, 14] scenarios. Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models [22, 9, 41, 45, 39], achieving superior scaling laws at low computational costs. Despite these advancements, MoE still faces efficiency challenges in both training and inference [3], and our work specifically focuses on enhancing inference-time efficiency.
@@
The imbalance in token-to-expert assignments [44, 5] poses a significant challenge to the deployment of MoE. This imbalance leads to inefficiencies in computation, communication, and memory [18, 37, 42], making it a critical bottleneck for MoE scalability and deployment. To mitigate this issue, an auxiliary balance loss [35] is incorporated into the training process to encourage more uniform token distribution across experts. Additionally, various training strategies have been introduced to further balance token assignments: Switch-Transformer [14] and DeepSeek-V2 [10] implement Token Drop to alleviate expert overload, while DeepSeek-V3 [11] introduces an additional sequence-level auxiliary loss to prevent severe token imbalance.
However, these techniques primarily focus on training and fail to ensure balanced token assignments during inference. Instead, addressing token imbalance at inference often incurs additional resource costs. For example, DeepSeek-V3 [11] mitigates this issue by duplicating high-load experts and deploying them redundantly. In contrast, our approach effectively balances token assignments without introducing additional computational overhead.
@@ 3.1 Extremely Imbalanced Expert Utilization
Figure 2 illustrates the normalized peak tokens load for each expert to accommodate all tokens within a single layer of OLMoE, where some experts receive an excessively large number of tokens (e.g., more than seven times the average number of tokens), leading to severe load imbalance and, consequently, significant latency. A detailed layer-by-layer analysis is provided in Appendix E.
@@ 4 Methodology
To address the question about overloaded experts, we first regulate their maximum utilization. Specifically, we introduce expert capacity to control token allocation. Given a capacity factor , the maximum number of tokens assigned to each expert (i.e., expert capacity) is defined as:
@@
where is typically much smaller than . This constraint ensures that no expert exceeds the specified capacity limit, effectively mitigating severe load imbalances and reducing latency. Note that tokens are distributed across devices under expert and data parallelism. To avoid additional communication overhead, we apply capacity constraints to tokens within each local device, similar to the constraints used during training [14]. This ensures that all devices respect the limits, maintaining strict control over token flow to the experts.
@@
where denotes the importance score of the mapping from the -th token to the -th expert. With this score, each overflowed expert selectively discards those with lower scores. Let be the set of overflowed experts:
@@
Order: Discarding later tokens once earlier tokens have filled the expert capacity. This strategy was first introduced in Switch-Transformer [14] during training, and we extend it to the inference phase.
@@
Score: Using the gating score as an importance indicator and discarding tokens.
Among these metrics, “Order” and “Reverse Order” are unstable, as shuffling sequences within a batch may result in different tokens being dropped [17]. “Random” assumes all tokens have an equal probability of being dropped. In contrast, “Score” is stable, unaffected by sequence order within a batch. Notably, there is virtually no additional computational overhead associated with calculating these metrics, and the dropping operation incurs minimal cost compared to the intensive computations performed by the experts.
@@ 5.1 Token Drop for High-load Experts
We next explore the efficiency improvements achieved by imposing expert capacity. Specifically, we employ distributed inference using eight H20 GPUs, utilizing an 8-way Data Parallelism (DP) and 8-way Expert Parallelism (EP) strategy through the Megatron-LM framework [36]. Notably, in Mixtral-87B-Instruct model, each GPU hosts a single expert, whereas, in models like OLMoE-Instruct, multiple experts must be deployed on a single GPU (e.g., eight experts per GPU) due to GPU resource constraints.
@@
Notably, the efficacy of acceleration is influenced by the numerical relationship between the total experts and the engaged GPUs in Expert Parallelism. As illustrated in Figure 6, for Mixtral-7B-Instruct, deploying a single expert per GPU maximizes the effectiveness of capacity-aware inference. In this configuration, Token Drop and Expanded Drop achieve end-to-end model speedups of 1.87 and 1.85, respectively, with . Conversely, deploying a greater number of experts on a single GPU results in more modest acceleration gains, as evidenced by the “8E/GPU” (OLMoE-Instruct and Deepseek-V2-Lite) the “10E/GPU” (Qwen1.5-MoE-Chat) in Figure 4 and Figure 6. This is because the aggregated load from multiple experts diminishes the proportion of reduced load, which is achieved by limiting the straggler expert. Therefore, it is anticipated that allocating more GPUs for expert distribution, thereby reducing the number of experts per GPU, would enhance the acceleration effect of capacity-aware inference.
@@
Figure 7 visualizes the number of dropped tokens across different capacity factors for various test datasets, with a more detailed illustration provided in Appendix F. Although the most overloaded expert receives much more tokens than the expected number of tokens , regulating the maximum capacity has a limited impact on the overall number of accommodated tokens, thereby maintaining competitive performance even after discarding overflow tokens. Moreover, dropping a small proportion of overflowed tokens can significantly reduce the latency caused by overloaded experts (e.g., dropping 12% of overloaded tokens promotes the inference speed by 85% in Mixtral-87B-Instruct), highlighting the efficacy of capacity-aware inference in improving both performance and efficiency.
@@
5.2 Expanded Drop to Low-load Experts
Besides the experts overloaded with tokens, some low-load experts receive only a few tokens, raising important questions: Are these low-load experts redundant and removable, or should they be leveraged to balance token allocation? Recent works [29, 19] remove less important experts to improve efficiency, while our proposed Expanded Drop increases their utilization by redistributing tokens for a more balanced assignment. Next, we investigate the significance of low-load experts and validate the effectiveness of Expanded Drop.
@@ 5.3 Extension to Multi-modal Mixture of Experts
In addition to applying capacity-aware inference to MoE models for language tasks, we also explore its effectiveness in multi-modal MoE settings. Specifically, we evaluate the OLMoE based MolmoE [12], across multi-modal benchmarks, including MME [15], MMBench [28], and SEED-Bench [27].
@@
Given the redundancy of image tokens and their large proportion in multi-modal tasks, we further investigate more aggressive capacity factors for Token Drop and Expanded Drop using the “Image First” strategy. Figure 10 demonstrates the effectiveness of Capacity-Aware Inference under low capacity constraints (i.e., ). This is largely due to the high redundancy in image tokens [4], which allows a higher dropping ratio without significantly affecting performance. Meanwhile, the dominance of image tokens enables the use of very low capacity factors without significantly affecting text token retention, as illustrated in Figure 10. Therefore, dropping image tokens at higher ratios leads to more balanced token assignments and substantially improved inference efficiency.
5.4 Ablation Study
Model-Specific Imbalanced Property
We explore the imbalance property in various models, such as OLMoE, DeepSeek, Qwen and Mixtral, which differ in both architectures (e.g., depth and width) and training strategies (e.g., training from scratch [31, 10] vs. training after upcycling [22, 39]).
On the one hand, our findings in Appendix E reveal different training strategies result in significantly varying levels of imbalance. Specifically, MoE models trained from scratch exhibit a much higher degree of imbalance. For instance, OLMoE and DeepSeek-V2-Lite experience peak expert-wise token allocations exceeding , whereas Qwen1.5-MoE and Mixtral are upcycled from dense language models, maintain a more balanced distribution, with peak expert-wise allocations staying below . This is because upcycling initializes all experts with identical parameters [24], reducing divergence and promoting balanced training in the early stages.
Capacity Factor

### v4

@@ Abstract
The Mixture of Experts (MoE) is an effective architecture for scaling large language models by leveraging sparse expert activation to balance performance and efficiency. However, under expert parallelism, MoE suffers from inference inefficiencies due to imbalanced token-to-expert assignment, where underloaded experts complete computations early but must wait for overloaded experts, leading to global delays. We define this phenomenon as the Straggler Effect, as the most burdened experts dictate the overall inference latency. To address this, we first propose Capacity-Aware Token Drop, which enforces expert capacity limits by discarding excess tokens from overloaded experts, effectively reducing load imbalance with minimal performance impact (e.g., speedup with only degradation on OLMoE). Next, given the presence of low-load experts remaining well below the capacity threshold, we introduce Capacity-Aware Expanded Drop, which allows tokens to include additional local experts in their candidate set before enforcing strict local capacity constraints, thereby improving load balance and enhancing the utilization of underused experts. Extensive experiments on both language and multimodal MoE models demonstrate the effectiveness of our approach, yielding substantial gains in expert utilization, model performance, and inference efficiency, e.g., applying Expanded Drop to Mixtral-87B-Instruct yields a 0.2% average performance improvement and a 1.85 inference speedup. The code is released at: https://github.com/CASE-Lab-UMD/Capacity-Aware-MoE.
@@ 1 Introduction
In recent years, the rapid evolution of Large Language Models (LLMs) (OpenAI, 2024; Team, 2024a; et al., 2024b) has driven a wave of innovations, continuously expanding the frontiers of AI research and applications. Among the model architectural innovations, the Mixture of Experts (MoE) framework has emerged as a pivotal technique for optimizing the cost-performance trade-off in LLMs. Specifically, MoE (Shazeer et al., 2017b) enhances scalability by integrating multiple experts while activating only a subset per input. This selective activation substantially improves model performance without a corresponding increase in computational cost, effectively balancing efficiency and performance.
Despite the success of MoE, a key efficiency challenge lies in the imbalanced token-to-expert distribution, which results in some experts being overloaded while others remain underutilized (Lepikhin et al., 2021; Zoph et al., 2022). In distributed GPU settings, experts are typically sharded across multiple devices, with each GPU responsible for a subset of the experts. Under expert parallelism, low-load experts complete their computations earlier but must wait for overloaded experts to finish, as synchronization barriers are required before proceeding to the next stage. This expert-level straggler effect further propagates to device-level delays, where GPUs hosting lighter expert workloads are stalled by GPUs hosting heavier workloads, leading to inefficient resource utilization and increased end-to-end latency during inference. As illustrated in Figure 1, this phenomenon is referred to as the Straggler Effect, where the heavily loaded experts determine the overall latency of imbalanced MoE inference.
While auxiliary balance losses have been incorporated into the training process to alleviate imbalance (Shazeer et al., 2017b; Fedus et al., 2022; et al., 2024b), these techniques remain ineffective in mitigating imbalance during inference. Specifically, as shown in Figure 1, our findings reveal a highly uneven token distribution among experts, with the highest-load expert handling more than seven times the expected average load. Moreover, managing such imbalance during inference often incurs additional resource overhead. For example, DeepSeek-V3 mitigates this issue by duplicating high-load experts and deploying them redundantly across devices (et al., 2024b). This motivates us to explore efficient token-to-expert assignment by addressing the key question: How can we prevent extreme overloading of heavily utilized experts?
@@
We explicitly clarify and analyze the Straggler Effect caused by test-time token imbalance in the Mixture of Experts, highlighting the optimization potential for reducing latency.
@@ 2 Related Works
The Mixture of Experts (MoE) is a kind of neural network architecture with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation (Jacobs et al., 1991; Jordan and Jacobs, 1994). The potential of sparse activation in MoE is subsequently exploited by Shazeer et al. (2017a) for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision (Riquelme et al., 2021) and language (Lepikhin et al., 2020; Du et al., 2022; Fedus et al., 2022) scenarios. Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models (Jiang et al., 2024; Dai et al., 2024; Xue et al., 2024a; Zhu et al., 2024; Team, 2024b), achieving superior scaling laws at low computational costs. Despite these advancements, MoE still faces efficiency challenges in both training and inference (Cai et al., 2024), and our work specifically focuses on enhancing inference-time efficiency.
@@
The imbalance in token-to-expert assignments (Zhou et al., 2022; Chen et al., 2022) poses a significant challenge to the deployment of MoE. This imbalance leads to inefficiencies in computation, communication, and memory (He et al., 2023; Song et al., 2023; Xue et al., 2024b), making it a critical bottleneck for MoE scalability and deployment. To mitigate this issue, an auxiliary balance loss (Shazeer et al., 2017b) is incorporated into the training process to encourage more uniform token distribution across experts. Additionally, various training strategies have been introduced to further balance token assignments: Switch-Transformer (Fedus et al., 2022) and DeepSeek-V2 (et al., 2024a) implement Token Drop to alleviate expert overload, while DeepSeek-V3 (et al., 2024b) introduces an additional sequence-level auxiliary loss to prevent severe token imbalance.
However, these techniques primarily focus on training and fail to ensure balanced token assignments during inference. Instead, addressing token imbalance at inference often incurs additional resource costs. For example, DeepSeek-V3 (et al., 2024b) mitigates this issue by duplicating high-load experts and deploying them redundantly. In contrast, our approach effectively balances token assignments without introducing additional computational overhead.
@@ 3.1 Extremely Imbalanced Expert Utilization
Figure 2 illustrates the normalized peak tokens load for each expert to accommodate all tokens within a single layer of OLMoE, where some experts receive an excessively large number of tokens (e.g., more than seven times the average number of tokens), leading to severe load imbalance and, consequently, significant latency. A detailed layer-by-layer analysis is provided in Appendix F.
@@ 4 Methodology
To address the question about overloaded experts, we first regulate their maximum utilization. Specifically, we introduce expert capacity to control token allocation. Given a capacity factor , the maximum number of tokens assigned to each expert (i.e., expert capacity) is defined as: .
@@
where is typically much smaller than . This constraint ensures that no expert exceeds the specified capacity limit, effectively mitigating severe load imbalances and reducing latency. Note that tokens are distributed across devices under expert and data parallelism. To avoid additional communication overhead, we apply capacity constraints to tokens within each local device, similar to the constraints used during training (Fedus et al., 2022). This ensures that all experts respect the limits, maintaining strict control over token flow to the experts.
@@
where denotes the importance score of the mapping from the -th token to the -th expert. With this score, each overflowed expert selectively discards those with lower scores. Let be the set of overflowed experts and the corresponding columns of , with the Token Drop threshold set as:
@@
Order: Discarding later tokens once earlier tokens have filled the expert capacity. This strategy was first introduced in Switch-Transformer (Fedus et al., 2022) during training, and we extend it to the inference phase.
@@
Score: Using gating scores after the softmax and top- operations as an importance indicator and discarding tokens.
Among these metrics, “Order” and “Reverse Order” are unstable, as shuffling sequences within a batch may result in different tokens being dropped (Hayes et al., 2024). “Random” assumes all tokens have an equal probability of being dropped. In contrast, “Score” is stable, unaffected by sequence order within a batch. Notably, there is virtually no additional computational overhead associated with calculating these metrics, and the dropping operation incurs minimal cost compared to the intensive computations performed by the experts.
@@ 5.1 Token Drop for High-load Experts
We next explore the efficiency improvements achieved by imposing expert capacity. Specifically, we employ distributed inference using eight H20 GPUs, utilizing an 8-way Data Parallelism (DP) and 8-way Expert Parallelism (EP) strategy through the Megatron-LM framework (Shoeybi et al., 2019). The input batches are configured with a batch size of 8K and a sequence length of 512, simulating real-time serving scenarios with high query throughput. Notably, in Mixtral-87B-Instruct model, each GPU typically hosts one or two experts, whereas, in models like OLMoE-Instruct, multiple experts must be deployed on a single GPU (e.g., eight experts per GPU) due to GPU resource constraints.
@@
Notably, the efficacy of acceleration is influenced by the numerical relationship between the total experts and the engaged GPUs in Expert Parallelism. As illustrated in Figure 6, for Mixtral-7B-Instruct, deploying one or two experts per GPU maximizes the effectiveness of capacity-aware inference. In this configuration, Token Drop and Expanded Drop achieve end-to-end model speedups of 1.87 and 1.85, respectively, with . Conversely, deploying a greater number of experts on a single GPU results in more modest acceleration gains, as evidenced by the “8E/GPU” (OLMoE-Instruct and Deepseek-V2-Lite) the “10E/GPU” (Qwen1.5-MoE-Chat) in Figure 4 and Figure 6. This is because the aggregated load from multiple experts diminishes the proportion of reduced load, which is achieved by limiting the straggler expert. Therefore, it is anticipated that allocating more GPUs for expert distribution, thereby reducing the number of experts per GPU, would enhance the acceleration effect of capacity-aware inference.
@@
Figure 7 visualizes the number of dropped tokens across different capacity factors for various test datasets, with a more detailed illustration provided in Appendix G. Although the most overloaded expert receives much more tokens than the expected number of tokens , regulating the maximum capacity has a limited impact on the overall number of accommodated tokens, thereby maintaining competitive performance even after discarding overflow tokens. Moreover, dropping a small proportion of overflowed tokens can significantly reduce the latency caused by overloaded experts (e.g., dropping 12% of overloaded tokens promotes the inference speed by 85% in Mixtral-87B-Instruct), highlighting the efficacy of capacity-aware inference in improving both performance and efficiency.
5.2 Expanded Drop to Low-load Experts
Some experts receive very few tokens, raising the question of whether they are redundant or can be leveraged for balanced allocation. We next examine their role and validate the effectiveness of Expanded Drop.
@@ 5.2 Expanded Drop to Low-load Experts
To analyze expert selection and justify Expanded Drop, we sort, for each token, all experts by their gating scores in descending order and record the ranked scores (top-1, 2, …, top-). Aggregating across tokens, we compute the average, maximum, and minimum score at each rank (Figure 8). The curves show that while top-ranked experts receive much higher scores, the decay across ranks is gradual rather than abrupt, yielding a relatively flat tail beyond the first few ranks. Consequently, experts just outside the top- often have scores comparable to those near the bottom of the top-, enabling rerouting or dropping without materially changing model behavior, i.e., rerouted tokens still go to reasonably relevant experts with similar gating probabilities. This validates the design of Expanded Drop, as it can exploit the flatness of the gating distribution to balance load without sacrificing accuracy.
@@
Advanced Variant: Device-Level Capacity-Aware Inference
In scenarios where multiple experts reside on the same device, the overall straggler effect is determined by the aggregated load on each device. Expert-level capacity constraints impose a strict limit. Specifically, the number of tokens assigned to a single expert must not exceed . When extended to a device with local experts, this implies that the total load is bounded by . However, this strict constraint can lead to over-aggressive token dropping, as a single overloaded expert may force token removal even when sufficient spare capacity remains across other experts on the same device.
To alleviate this issue, we introduce a device-level capacity-aware formulation, which applies the constraint at the device granularity rather than on individual experts. For instance, we enforce . Unlike expert-level constraints, which may drop tokens simply because a single expert exceeds its local budget, the device-level constraint admits these tokens as long as the total device-level load remains within the allowable bound. This results in a smoother, less restrictive, and more utilization-efficient alternative.
| Method | Granularity | OBQA | PIQA | RTE | WinoGrande | BoolQ | ARC-C | HellaSwag | MMLU | GSM8K | Avg. | |
| Baseline | – | 45.2 | 80.6 | 82.3 | 69.5 | 88.6 | 69.5 | 77.6 | 77.9 | 89.5 | 75.6 | |
| Expanded Drop | Expert | 43.6 | 79.4 | 80.3 | 68.4 | 86.5 | 68.8 | 74.4 | 76.8 | 87.0 | 73.9 | |
| Device | 42.4 | 79.7 | 81.2 | 69.3 | 88.0 | 69.4 | 77.1 | 77.1 | 88.9 | 74.8 |
We evaluate this variant and observe that device-level constraints consistently outperform expert-level constraints in downstream performance, as shown in Table 3. This more flexible constraint also leads to stronger practical speedups: on Qwen3-MoE (Yang et al., 2025) with , the device-level formulation achieves a end-to-end speedup and a speedup on a single MoE layer, both surpassing the and speedups obtained under expert-level constraints with .
@@ 5.3 Extension to Multi-modal Mixture of Experts
In addition to applying capacity-aware inference to MoE models for language tasks, we also explore its effectiveness in multi-modal MoE settings. Specifically, we evaluate the OLMoE based MolmoE (Deitke et al., 2024), across multi-modal benchmarks, including MME (Fu et al., 2023), MMBench (Liu et al., 2023), and SEED-Bench (Li et al., 2024a).
@@
Given the redundancy of image tokens and their large proportion in multi-modal tasks, we further investigate more aggressive capacity factors for Token Drop and Expanded Drop using the “Image First” strategy. Figure 11 demonstrates the effectiveness of Capacity-Aware Inference under low capacity constraints (i.e., ). This is largely due to the high redundancy in image tokens (Chen et al., 2024), which allows a higher dropping ratio without significantly affecting performance. The dominance of image tokens also enables the use of very low capacity factors without significantly affecting text token retention, as illustrated in Figure 11. Dropping image tokens at higher ratios leads to more balanced token assignments and substantially improved inference efficiency.
@@ 6 Conclusion
7 Ethics Statement
Our work focuses on improving the efficiency of large-scale Mixture-of-Experts (MoE) models at inference time. The proposed methods are lightweight modifications to model execution and do not involve collecting or annotating new data. All experiments are conducted on publicly available pretrained models and standard open-source benchmarks. We emphasize that our contributions are intended solely for research and educational purposes in efficient machine learning. We will explicitly release code under a research license and include clear terms discouraging misuse in sensitive applications such as mass surveillance or privacy-intrusive deployments.
8 Reproducibility Statement
We are committed to ensuring full reproducibility of our results. Upon acceptance, we will release the complete implementation of Capacity-Aware Inference, including inference-time strategies (Token Drop and Expanded Drop), evaluation harness integration, and scripts for running all benchmarks. Detailed instructions will be provided in a reproducibility README covering environment setup, library versions, random seeds, and commands for end-to-end replication.